# Teaching Divisibility and Binomials with Coq

### v1

@@ 3.2.2 Worksheet
In order to help students, we may want to reduce what the Search command gives them, and this is done in the WS_Helper module. For the following lemma, we only need
Export Nat (mul_1_r, pow_1_l).
And finally the lemma to be shown is stated. We can also imagine a version where the Coq statement is written by the student from the mathematical one: $\forall n,\in\mathbb{N},\sum_{k=0}^{n}{n\choose k}=2^{n}$.
@@
(*fr Transformer 2 en une somme. *) (*en Transform 2 into a sum. *)
(*fr Utiliser la formule du binome. *) (*en Use binomial theorem. *)
(*fr Reduire l’egalite de somme a l’egalite des fonctions. *)
(*en Reduce equality on sums to equality on functions. *)
@@
replace 2 with (1+1) by reflexivity. replace 2 with (1+1) by easy.
rewrite binom_formula. rewrite binom_formula.
apply sum_range_ext. intros k Hk. apply sum_range_ext. intros k Hk.
rewrite !pow_1_l. rewrite !pow_1_l; lia.
rewrite !mul_1_r. reflexivity. Qed.
Qed. (* End other solution for the teacher. *)
@@
A script automatically generates the student version with comments in the selected language. It replaces the tagged solution with command Admitted so that the file is still compiling. Regarding the automation, there are several possible levels, as exhibited by the previous example. Another question is the place of variables in statements. In Coq, we may write Lemma name : $\forall$ n, or Lemma name n : Both scripts are equivalent, and have their pedagogical advantages and disadvantages that are not discussed here. We have made an arbitrary choice that can be modified by the teacher.
@@ 4 Conclusion
We have shown a Coq development about divisibility and binomials. It consists in additional lemmas, dedicated tactics and 43 exercises in two worksheets. The total is more than 250 lemmas including the 43 exercises, and is 2,500 lines of Coq long, including about 1,000 lines of specification and comments. Ideas of exercises either come from textbooks [2, 1] or the internet, mostly bibmath.net; a few directly come from a worksheet of a teacher.
@@
This library is free. Teachers may use part or all of it, with several levels of leeway as explained above: which exercises (from the worksheet or even the library), which automation, which environment. This pedagogical freedom is important to adapt the contents to (i) the level of the students and (ii) the goal of the course. We focus on the early academic years or even a year before, but it means a span of four years leading to more or less mature students. The goal may be either purely mathematics, or mainly logic or a trade-off between both. For example, a teacher may want to prove more advanced lemmas with as much automation as possible while another may prevent automation to focus on deduction rules and logical quantifiers.
@@
A difficult point is how this work relates with the Coq standard library, as we need some of its lemmas but would like to replace others. Another difficult point is the level of notation and abstraction allowed. It is tempting to abstract mathematical objects and make the proofs once, but it creates strange hypotheses as seen in Section 2.1. If a student tries to unfold all definitions, both abstraction and notation may make the goal impossible to read and understand.
@@
The scope of this work is only integer properties. But we think we have spanned over many types of exercises, as done in mathematics. We have studied the ones we were not satisfied by the proof in order to develop either additional lemmas or tactics to make them reachable. The contribution is therefore both a Coq library and several Coq worksheets, done with a focus on flexibility and a pedagogical insight. A first perspective is of course for this material to be used on real students, be it by ourselves or not. It may help sharpen some of the exercises. Our next step is more advanced mathematics and we plan to develop a worksheet on the Lebesgue integral. This is at the other end of the early academic years as possible (usually done at the end of the bachelor, three years after terminale). It deals with advanced concepts such as real numbers, functions, convergence. The mathematics usually rely on implicit axioms such as extensionality or classical logic. We plan to use a Coq library developed for purely research purpose [3] and see how it can be adapted to obtain pedagogical contents.
@@ Acknowledgments
We are thankful to Mr. Delabrouille for letting us use his paper worksheet exercises for divisibility.

### v4

@@ 3.3 Worksheet
Proof. Proof.
(*fr1 Indication : essayer n = 89. *) (*fr2 Indication : trouver une valeur pour n *)
(*en1 Hint: try n = 89. *) (* en utilisant un papier et un crayon. *)
(*en2 Hint: find a value for n *)
(* by using pen and paper. *)
(* Begin solution for the teacher. *)
exists 89. split.
+ apply sub_0_le. rewrite 5!sub_succ. apply sub_0_l.
+ apply (mul_cancel_l _ _ 5). discriminate.
rewrite binom_mul_S_r.
rewrite 4!sub_succ, sub_0_r, mul_assoc.
reflexivity.
Qed.
(* End solution for the teacher. *)
For this kind of basic exercise, it could be more convenient to deal with the factorial expression (1) with the division of real numbers (instead of the Euclidean division). But this would add the difficulty to deal with the injection from $\mathbb{N}$ to $\mathbb{R}$.
The second example, binom_sum_sqr (from the library), illustrates the possibility to give the teacher several proofs, here with automation (on the right) or without (on the left). We have chosen a mild level of automation, where only trivial notions with no consequence on the mathematical path are proved with automatic tactics (here lia and now). We can also imagine a version where the Coq statement is written by the student from the mathematical one (11).
@@
(*fr Indications :
- Transformer (2*n) en une somme.
- Utiliser l’identité de Vandermonde.
- Réduire l’égalité de somme à  l’égalité des fonctions.
- Utiliser la symétrie des coefficients binomiaux.
- Terminer avec les conditions d’application des lemmes utilisés. *)
(*en Hints:
- Transform (2*n) into a sum.
- Use Vandermonde’s identity.
- Reduce equality on sums to equality on functions.
- Use symmetry of binomial coefficients.
- End with application conditions of used lemmas. *)
@@
replace (2 * n) with (n + n) replace (2 * n) with (n + n) by lia.
by apply double_twice. rewrite binom_vandermonde.
rewrite binom_vandermonde. apply sum_range_ext. intros k Hk.
[…]
2: assumption. Qed.
rewrite pow_2_r. reflexivity. (* End other solution for the teacher. *)
Qed.
@@
We can note that the WS_Helper whitelist mentioned above can reduce here to:
Export Nat (double_twice, le_add_r, lt_succ_r, pow_2_r).
Additional hint in the worksheet could be the output of a Coq command, for instance:
(* Check binom_vandermonde:
[…]
The third example, binom_hockey_stick (also from the library), illustrates the possibility to give teachers three levels of hints (from most to least detailed) on a more elaborate property.
(*fr Formule de la gouttière. *) (*en Hockey stick identity. *)
(*fr Faire la preuve du lemme suivant et remplacer "Admitted" par "Qed". *)
(*en Fill the proof of next lemma and replace "Admitted" by "Qed". *)
[…]
Proof.
(*fr1 Indications :
- Commencer par énoncer une version équivalente en remplaçant "sum_range k n" par
"sum_n k p", où p représente le nombre de termes sommés, c’est-à -dire n-k+1.
Donc "binom (S n) (S k)" devient "binom (k + p) (S k)" et il faut prouver
[…]
- Prouver cette version par induction sur p, en utilisant la formule de Pascal pour
le cas successeur.
- Puis faire apparaître "sum_n", utiliser le résultat prouvé précédemment, et
discuter suivant que "S n < k" ou pas. *)
(*en1 Hints:
- First state an equivalent version by replacing "sum_range k n" by "sum_n k p",
where p stands for the number of added terms, ie n-k+1.
Thus "binom (S n) (S k)" becomes "binom (k + p) (S k)" and we have to prove
[…]
- Prove this version by induction on p. Use Pascal’s rule for the successor case.
- Then, make "sum_n" visible, use the result proved previously, and discuss
whether "S n < k" or not. *)
(*fr2 Indications :
- Commencer par énoncer une version équivalente en remplaçant "sum_range k n" par
"sum_n k p", où p représente le nombre de termes sommés, c’est-à -dire n-k+1,
donc "binom (S n) (S k)" devient "binom (k + p) (S k)".
- Prouver cette version par induction sur p (penser à  la formule de Pascal).
- Utiliser le résultat prouvé précédemment, et discuter suivant que "S n < k" ou pas. *)
(*en2 Hints:
- First state an equivalent version by replacing "sum_range k n" by "sum_n k p",
where p stands for the number of added terms, ie n-k+1, thus "binom (S n) (S k)"
becomes "binom (k + p) (S k)".
- Prove this version by induction on p (think about Pascal’s rule).
- Then, use the result proved previously, and discuss whether "S n < k" or not. *)
(*fr3 Indications :
- Commencer par énoncer une version équivalente en remplaçant "sum_range k n" par
"sum_n k p", où p représente le nombre de termes sommés.
- Prouver cette version par induction sur p, puis utiliser ce résultat intermédiaire. *)
(*en3 Hints:
- First state an equivalent version by replacing "sum_range k n" by "sum_n k p",
where p stands for the number of added terms.
- Prove this version by induction on p, then use this intermediary result. *)
(* Begin solution for the teacher. *)
[…]
intros p. induction p.
- intros. simpl. rewrite binom_eq_0.
reflexivity.
rewrite add_0_r. apply lt_succ_diag_r.
- intros. simpl. rewrite IHp.
[…]
- unfold sum_range; rewrite H. destruct (nat_lt_le_dec (S n) k) as [H1 | H1].
+ rewrite !binom_eq_0. (* lia solves next subgoals. *)
reflexivity.
apply lt_trans with k. assumption. apply lt_succ_diag_r.
rewrite (nat_sub_0_lt _ _ H1). rewrite add_0_r. apply lt_succ_diag_r.
+ f_equal. (* lia solves the next subgoal. *)
rewrite (add_sub_assoc _ _ _ H1).
rewrite add_comm, add_sub. reflexivity.
Qed.
(* End solution for the teacher. *)
Another question is the place of variables in statements. In Coq, we may write
Lemma name : $\forall$ n,     or     Lemma name n :
Both scripts are equivalent, and have their pedagogical advantages and disadvantages that are not discussed here. We have made an arbitrary choice that can be modified by the teacher.
4 Experiment
In June 2024, some of us co-supervised, with Christine Paulin, two first-year undergraduate students during two weeks. They come from a dual degree curriculum, and had already completed a course on formal methods using Lean and a waterproof-like library as support. They expressed interest in an internship related to formal methods so we took this opportunity to let them test our worksheets.
They did not have any experience using Coq. We installed CoqIde on their machine, and let them finish an introductory worksheet [15] to get to know Coq. In less than two weeks, they managed to finish the two worksheets we had created. They were very autonomous for first-year undergraduate students.
We took into account several unsatisfactory points:
  • The students relied on the Check command to interpret what they were seeing in the goal view. This is generally fine, except when coercions are involved. They did not understand how statements using coercions could possibly typecheck, and were generally very confused about them. We therefore removed the coercions (see Section 2.4), even though it made the statements longer. The students preferred this, and were able to complete the exercises.
  • Another command the students used was Search. While they preferred using it rather than searching manually, they complained that it showed too much theorems that were unrelated. Our solution to this problem was to use the Add Search Blacklist command. This removes from the search result all theorems from the standard library. Then, we made our own module Z, which contains only theorems and definitions we want the students to use. This solves the issue, and Search now only shows relevant definitions and lemmas. This is especially the case for binomials, where the line is blurred between a (standard) library and some of our exercises.
Another interesting point is that, for the exercise binom_committee_chair_alt, they managed to complete the proof using another longer way we had not thought about.
@@ 5 Conclusion
We have shown a Coq development about divisibility and binomials. It consists in additional lemmas, dedicated tactics and 49 exercises in two worksheets. The total is more than 250 lemmas including the 49 exercises, and is 3,000 lines of Coq long, including about 1,300 lines of specification and comments. Ideas of exercises either come from textbooks [3, 2] or the internet, mostly bibmath.net; a few directly come from a worksheet of a teacher.
@@
This library is free. Teachers may use part or all of it, with several levels of leeway as explained above: which exercises (from the worksheet or even the library), which automation, which environment, which level of hints. This pedagogical freedom is important to adapt the contents to (i) the level of the students and (ii) the goal of the course. We focus on the early academic years or even a year before, but it means a span of four years leading to more or less mature students. The goal may be either purely mathematics, or mainly logic or a trade-off between both. For example, a teacher may want to prove more advanced lemmas with as much automation as possible while another may prevent automation to focus on deduction rules and logical quantifiers.
@@
A difficult point is how this work relates with the Coq standard library. It may be inconvenient to require the standard library with its bunch of lemmas that are probably too much for a beginning student (in a Search for instance). But it is inconvenient too to not require it and have a parallel development, to be maintained too. Another difficult point is the level of notation and abstraction allowed. It is tempting to abstract mathematical objects and make the proofs once, but it creates strange hypotheses as seen in Section 2.1. If a student tries to unfold all definitions, both abstraction and notation may make the goal impossible to read and understand.
@@
The scope of this work is only integer properties. But we think we have spanned over many types of exercises, as done in mathematics. We have studied the ones for which we were not satisfied by the proof in order to develop either additional lemmas or tactics to make them reachable. The contribution is therefore both a Coq library and several Coq worksheets, done with a focus on flexibility and a pedagogical insight.
Perspectives
A first perspective is of course for this material to be used more intensively on real students. It may help sharpen some of the exercises, by their hints, stating or expected difficulty.
A second perspective is about the script used to generate the version for students. By now, it is designed to keep only one language and it would be interesting to keep at least two of them. For instance, when selecting the local language, it would be interesting to also keep English, e.g. for the foreign students. Moreover, as presented with the worksheet about binomials, it will be interesting to handle the possibility to provide several levels of indications, e.g. with the tags fr1/en1, fr2/en2 and so on.
Another perspective is more advanced mathematics and we plan to develop a worksheet on the Lebesgue integral. This is at the other end of the early academic years as possible (usually done at the end of the bachelor, three years after terminale). It deals with advanced concepts such as real numbers, functions, convergence. The mathematics usually rely on implicit axioms such as extensionality or classical logic. We plan to use a Coq library developed for purely research purpose [4] and see how it can be adapted to obtain pedagogical contents.
@@ Acknowledgments
We are very thankful to Gilles Delabrouille for letting us use his paper worksheet exercises for divisibility.
We are also very thankful to Julie Delafosse and Noé Briand for testing our worksheets during their internship. Their insights and efficiency have impressed us. Finally, we are very thankful to Christine Paulin for co-supervising the two students of the experiment.